### v1

# Affordances and Design Principles of The Political Left and Right

Abstract.
Like any form of technology, social media services embed values. To examine how societal values may be present in these systems, we focus on exploring political ideology as a value system. We organised four co-design workshops with political representatives from five major parties in Finland to investigate what values they would incorporate into social media services. The participants were divided into one right-leaning group, two left-leaning groups, and one mixed group. This approach allows us to examine how their political ideologies, i.e., value systems, influenced the design of social media. We analysed produced artefacts (early-stage paper mockups) to identify different features and affordances for each group and then contrasted the ideological compositions. Our results revealed a clear distinction between groups: right-leaning groups favoured market-based visibility, while left-leaning groups rejected such design principles in favour of open profile work. Additionally, we found tentative differences in design outcomes along the liberal–conservative dimension. These findings underscore the importance of acknowledging existing political value systems in the design of social computing systems. They also highlight the need for further research to map out political ideologies in technology design.
political ideologies; politicians; left—right ideology; liberal–conservative ideology; co-design
copyright: $c$
1. Introduction
Technology design always encompasses values, a well-known fact among human-computer interaction researchers and practitioners. To account for various values, value-sensitive design approaches were developed to help align the system’s values with its user’s values (Friedman et al., 2009; Borning and Muller, 2012). Extensive work has been done to understand the values of children and parents in the context of parental software (Van Mechelen et al., 2014; Nouwen et al., 2015). Alternatively, some design work incorporates specific values as a political statement, highlighting a social issue or envisioning alternative futures. For example, Fox et al. (2019) focused on the distribution of menstrual products in public toilets, a feminist issue.
Researchers have often focused on end-users and less on decision-makers, such as political elites, when considering different values and their influence on design. At the same time, the government is a major provider of services, facilitated increasingly through digital means. Even more importantly, in democratic societies, the elected representatives have the legislative power – they establish laws increasingly related to technology, as shown in the EU AI Act or the California Consumer Privacy Act. Both involve value-based decisions about the rights and responsibilities of citizens and their expected roles. In the political system, there are reasons to suggest that such decisions mirror a wider set of beliefs and assumptions (or political ideologies), as well as political realities based on different balances of power among representatives. Members of Congress, Parliament or city councils, and other representatives rarely have been studied as stakeholders in the design of digital systems.
To this end, we organised co-design workshops with political representatives to elicit how their ideologies would be embedded in technology design. In four co-design workshops, 13 participants across the political spectrum were tasked to design a social media platform. We arranged participants into one right-leaning group, two left-leaning groups, and one mixed group for the workshops to compare their design and value decisions.
We examined the outcomes of the workshops as the participants’ suggestions and analysed those towards the specific features and implicit affordances. Based on the literature on social media platforms (Boyd, 2010; Bucher and Helmond, 2017; DeVito et al., 2017; Van Raemdonck and Pierson, 2021), we categorised affordances and connected them to underlying values to uncover ideological tensions between the different design solutions.
We next discuss relevant literature to present political ideologies as values and discuss social media design further. Following this, we describe the organisation of the co-design workshops. In this paper, we focus only on the final artefacts of the co-design activity, which exemplify the value-based decisions the groups made during the design workshop. We present a method to analyse these artefacts and summarise observations from the artefacts. We observe that right-leaning groups supported market-based dynamics while left-leaning groups supported diverse self-presentation. Mixed groups seemed to have both kinds of features. Before concluding our work, we discuss the importance of ideological diversity in design activities. Our work makes the following contributions focused on the design of information technology:
  • We introduce political ideologies as a set of values and connect the debate on values to national politics.
  • We introduce a co-design approach which allows one to examine political ideologies in technology design.
  • We show how political ideologies (right-left) shape social media design and identify design principles of flexible identity and competitive logic to be specific to the left- and right-leaning parties.
2. Background
Our work is based on the well-established observation that technology design and development are not value-neutral activities (among many others Nissenbaum, 2005; Winner, 1985). Winner (1985) infamously asked if artefacts have politics following a story about the bridges on Long Island, New York, USA. The bridges were constructed too low for public transportation buses to pass under them, thus removing access to the area for people using public transportation, primarily black people at that time. He suggests that this is related to the designers’ social class bias and racial prejudice. Therefore, the bridges became manifestations of their designer’s political values. In the context of human–computer interaction, Keyes et al. (2019) has similarly called out the often apolitical nature of work, which they say to be implicitly neo-liberal in nature, and thus calling out for more diverse ideological work.
2.1. Political Ideologies as a Set of Values
Across various disciplines, there are several ways to conceptualise values. Political scientists often focus on political ideologies instead of values. Political ideologies are a set of values, beliefs, and attitudes people share that impact how they perceive and act in the world (Denzau and North, 1994). Instead of single values, they formulate a value system. Within representative bodies, people with similar ideologies form a political party (or a social movement in their early stages) to collaborate in advocating their values. The existence of cohesive parties allows one to map out ideologies’ dimensions; in the United States of America, this is often represented as a single dimension (republican–democrat) where various societal conflicts, such as economic stances, the role of the government, or racial assumptions, are positioned.
Figure 1. The two-dimensional model of political ideologies
[…]
Two axes crossing with one labelled value liberal and value conservative, and the other labelled economic right and economic right.
In Europe (where our study is situated), the multiparty system makes more differentiated ideologies apparent. Instead of a single axis, a two-dimensional model is used across these countries (see Figure 1). The first axis corresponds to different values in terms of the degree of economic freedom (left–right), and the second axis corresponds to the level of desirable societal diversity (conservative–liberal or green-alternative-liberal–traditional-authoritarian-nationalist) (CAUGHEY et al., 2019; Kriesi et al., 2006). For example, a supporter or member of a left-liberal party, such as social democrats or the greens, might push for the right to self-define gender identity and extensive government oversight in the economic operation. In contrast, a supporter or member of a right-conservative party, such as Christian democrats or conservative parties, might push for traditional gender roles and limit the government’s role in the economy. In general, while the model is simplistic, it has provided extensive value for empirical research on political decision-making, political opinions, and voting (e.g., CAUGHEY et al., 2019; Huber and Inglehart, 1995; Hooghe et al., 2002; Knight, 2006). There is mixed evidence that left-leaning political parties are adapting information and communication technology more extensively to allow contact and participation in party functions (González-Cacheda and Cancela Outeda, 2024).
As ideologies are systems based on a set of values, these values can manifest themselves in the design and use of digital services. Examining the design of social media services, Grön and Nelimarkka (2020) illustrate that the left–right axis relates to different degrees of paternalistic values: left-leaning participants supported a higher degree of paternalism, while right-leaning participants highlighted the importance of individual responsibility. Their work highlights ideologies as an area of political competition, where different ideologies prefer design approaches aligned with their worldview. Similarly, Détienne et al. (2019) built on the concept of ideologies to understand community memberships in conflict. For example, they identified an ideological conflict in the editing of Wikipedia articles. They observed strong disagreement between authors with a religious point of view and those with a scientific point of view, leading to an edit war and personal attacks on the platform. This instance underscores the potential for conflict and disagreement between political ideologies, making it clear that finding a compromise on an agreed solution can be difficult and sometimes impossible. This work highlights the importance of reflecting more expansive sets of values (or ideologies) and understanding them through wider groups, even if their ideologies do not fit the two-dimensional political ideology model presented above.
Although ideologies and their impact on information systems design have received less attention in human–computer interaction, the above examples illustrate the value of approaching not only individual values but sets of values. The focus on political ideologies – pushed forward by political parties – provides the additional benefit of reflecting the legislative and executive processes at the societal level.
2.2. Values in the Context of Social Media Services
Like any socio-technical arrangement, social media services incorporate and display values. The most prominent value tension in contemporary public discourse concerns social media companies’ for-profit operations and their contrast with other values, such as human rights. Thus, the discourse has a focus on capitalism. For example, Zuboff (2018) has argued that we are now in the era of surveillance capitalism, in which the collection and use of personal data has become a widespread business model. Social media companies use such personal data for targeted advertising, but while doing so, limit our right to privacy. An alternative line of criticism is the focus on user engagement and attention as a (capitalist) metric, leading to the promotion of affective content and thus having negative implications for political discourse on these platforms.
Given these concerns, there are proposals on alternative value bases for social media services. Jia et al. (2024) discuss how the news-feed algorithm could promote content more aligned with democratic ideals by sorting content so that it does not show support for undemocratic practices, partisan animosity or violence, or social distrust. Others have proposed different normative bases for rethinking social media services (for summary, see Goldberg, 2016). For example, the ideas that social media services could serve as a place for political deliberation have been repeated in the literature (Semaan et al., 2014, 2015, among others). These examples again illustrate the critical role of values as a standing point for social media service design.
However, beyond the above-mentioned Grön and Nelimarkka (2020), social media design research has not examined values through the lens of political ideologies. Instead, it seems that political ideologies form a design challenge in the context of social media service design: how to ensure that people across the ideological spectrum interact to create common ground (among many others Combs et al., 2023; Feltwell et al., 2020; Garimella et al., 2017). Alternatively, the broader social media research sees ideology as a background variable explaining human behaviour or content produced (for example, Morgan et al., 2013; Zhang and Counts, 2015).
Our brief review highlights the role of values in the debate on social media service design. It critiques the current situation and suggests alternative design approaches. At the same time, we highlight that, from this perspective and more widely in human–computer interaction research, these values are rarely understood in terms of political ideologies.
3. Approach, Data and Methods
Our design activity focused on social media services as these are increasingly critical for information dissemination and political discourse in our society. Our form of co-design aligns with “co-creative design”, where the collaborative creativity of participants and designers creates value, which, in our case, are insights towards the research inquiry. Here, participants are not automatically stakeholders (Zamenopoulos and Alexiou, 2018; Prahalad and Ramaswamy, 2004).
3.1. Research Setting
This study is situated in Finland, a Nordic country with 17 registered political parties, of which nine have representatives in the Finnish Parliament. Like in the rest of Europe, the media and political scientists use the two-dimensional ideological model (Figure 1) to categorise these parties (Grönlund and Söderlund, [n. d.]). In recent years, the ideological dimensions have become more prominent and affective polarisation has increased (Kekkonen et al., 2024).
In terms of political power, there are three equally large parties, each with about 20% of the popular vote in the most recent elections: National Coalition Party, the Finns Party, and Social Democratic Party. National Coalition Party and the Finns Party are political right, whereas the Social Democratic Party is political left. The Finns party is the only conservative party in this group. Following them, there are three somewhat smaller parties, with between 5% and 10% of the popular vote: the Center Party, the Greens Party, and the Left Alliance, The Greens party and the Left Alliance are left-leaning parties, while the Centre party is right-leaning and the only conservative party in this set of parties. All parties above have elected municipal, national and European representatives and have been part of the current or previous national government. The three remaining parties Swedish People’s Party of Finland Christian Democrats, and Movement Now, all had less than five per cent of the popular vote.
3.2. Participants
Figure 2. Position of workshop participant on the ideological map.
[…]
In total, we organised four workshops with 13 participants representing five111 We reached out to all parties, but even after several attempts, one party did not respond to our request. This also limited our ability to organise mixed and right-leaning workshops. of the six major parties. These parties included both right-leaning and left-leaning parties and included parties with more centralist stances and more extream stances. These parties represent the Finnish political spectrum broadly: together these five parties gained more than 60% of the popular vote, with the missing party adding an additional 20% of popular vote to the mix; thus our original target represented 86% of total popular vote. We contacted the party’s general secretary, executive director, or similar high-level officials. We asked them to suggest candidates for the workshop with expertise in social media and familiarity with the party’s political positions. Their suggestions included city council members, party office staff and youth organisation members, who we would consider to have a clear political ideology. Of the participants, six identified as women, six identified as men, and one identified in the other gender category. Their ages ranged from the early twenties to the seventies, with a median of 34 and a mean of 39. As Finland is a relatively small country, we do not provide more detailed demographic details to ensure the anonymity of participants.
The candidates suggested by party officials were suitable participants for this study. First, they have experience discussing politics through computer-mediated means and face-to-face settings. All participants discussed politics daily or often in face-to-face settings. Majority also discussed politics daily or often through instant messaging apps (5 daily, 6 often and 2 sometimes) and on social media (7 daily, 2 often and 4 sometimes). Second, they had an established ideological stance. All participants were able to indicate a closer political party and this party position was well established.
The groups had varied ideological compositions (see Table 1): one consisted of solely right-leaning participants (Group 1), two consisted of solely left-leaning participants (Groups 2 and 3), and one were mixed, including both right and left-leaning participants (Group 4). With the exception of Group 1222 Disappointingly, this group had a now-show from the other right-leaning party in our participant pool. This lead to the unfortunate situation, that this group was homogeneous in terms of political parties. As the logic of the workshops was complicated, we opted to run this group to ensure that we have at least one right-leaning group in our sample., all groups had participants from several parties. This group composition allowed us to examine the elicitation of political preferences in the left–right leaning axis, thus engaging how ideological thinking impacts research activities.
Group 1 Group 2 Group 3 Group 4
Right-leaning Left-leaning Left-leaning Mixed
Left (0)–Right (10) 8.0 1.33 3.50 5.50
(0.00) (0.577) (1.29) (3.00)
Conservative (0)–Liberal (10) 6.00 7.00 8.75 6.75
(2.83) (3.46) (0.96) (3.40)
Table 1. Mean and standard deviation of ideological measurements per group.
3.3. Organisation of Co-design Workshops
The co-design workshop lasted about 2–3 hours and consisted of five activities that together (see Table 2). The first two activities were individual orientation and thinking activities. First, participants were asked to atune to the problems of social media and after, they were asked to experiment with the materiality of the social media service by considering its issues and challenges through primarly through an user interface, tasking users to consider its modifications. These activities were designed as individual to reduce group-think and peer pressure, allowing participants to form their own perspectives. Following the two individual activities, the participants were asked to share their thoughts and observations with the group and briefly introduce themselves if they had not done so already. This activity allowed all participants to share perspectives and observations without interruption. Following this, the participants moved on to the co-design activity, where they were assigned to design a new social media service in about 45–60 minutes. We outline the problem statement around political polarisation and decreasing political discussion culture, with the following brief:
Activity Assignment Minutes
Individual orientation activity The political debate on social media presents a range of opportunities and challenges. What would you highlight as challenges and opportunities? Participants were instructed to use sticky notes during, with one idea per sticky notes. 10
Individual design thinking activity What Twitter as a service could do to reduce the challenges and highlight the better opportunities. Think about this through the different perspectives of Twitter: what should change? Participants were given mock-up illustrations from Twitter and instructed to make notes and draws over them. 10–15
Sharing activity Please share key observations and highlights to the group. 15
Co-design activity Design a new social media service to address issues on social media you have observed. (More extensive brief presented inline.) 45–60
Presentation and survey Please present the prototype to the research team and answer their questions. Please fill in an exit survey. 15–25
Table 2. Overview of the co-design workshop schedule
Your task is to work together to design a new service for people who want to discuss politics: PoliPulse. The service aims at politically active Finns to exchange opinions and reduce polarisation in society. The aim is to avoid the harmful features of other social media and to improve the functionality and atmosphere of social debate. You can only post text and links to the service. Your task as a group is to create an initial prototype of this service, showing the main interfaces and being able to explain what users do in these interfaces.
During the workshop, we did not want to intervene with the group dynamics and reasoning processes, so the research team did not facilitate the co-design activity beyond providing assignments and time management. The exception was ensuring that all participants had their say during the sharing activity if the groups were not doing this themselves. During the co-design activity, we provided the groups with pre-made icon sets and neutral figures to reduce the need for graphic design skills; they were also given empty user interface canvases and pens. We did not elaborate on their interpretations of the assignments or design briefs beyond practical considerations – such as reminding people to use one sticky note per idea, use legible handwriting, or introduce the design materials.
In the final step, each group needed to present their joint proposal, which demanded that groups consolidate which ideas to include in their final artefact. We were primarily interested in these artefacts as the basis for our analysis, as they should exemplify ideological tendencies in the outcomes of a design process.
3.4. Ethical Considerations
Our methods borrow from the long-standing traditions of participatory design, so we want to reflect on our use of collaborative design as a research method. In our case, we purposely gathered members of the political elite instead of any under-represented group, aligning with the tradition of ‘studying up’. As Nader (1972) has highlighted, focusing research on those in positions of power opens up new ways to theorise and informs democratic societies on the role of the powerful. As we introduce novelty in using co-design workshops to study connections between politics and design, we caution other researchers to use these with respect to the political realities in place.
3.5. Analysis
The primary aim of this work was to highlight how political ideologies become embedded in the design process. The workshop we conducted served to gather the cultural norms, assumptions, and ideologies of the co-designers, here the political representatives, by capturing a design process. Co-design activities include examining various design directions and ideas; however, for this work, we wanted to focus on each group’s converged presentation. Therefore, we focus only on the co-designed artefacts, i.e. sketches on potential user interfaces (see Figure 3 for an example).
We used qualitative content analysis to identify affordances from these sketches (Rodgers et al., 2000). We build on the Feature Analysis method (Hasinoff and Bivens, 2021) to connect this analysis to ideologies to examine design artefacts for their value-loaded contents. Hasinoff and Bivens (2021) developed the Feature Analysis as a qualitative content analysis that identifies ideologies in mobile applications based on their features. Their work builds on an inductive approach to identifying features (and later affordances) and grouping them to examine ideologies. As we already have ideological focuses of groups, we focus more on examining if there are patterns in the design approaches.
Naturally, the question becomes what is a feature or an affordance. For Hasinoff and Bivens (2021), a feature was “a function that users control or are likely aware of”. In contrast, an affordance was a higher level conceptualisation of a characteristic of the technology that “request, demand, allow, encourage, discourage, and refuse” (Davis and Chouinard, 2016, p. 241) a user behaviour. As affordances have been widely used in human–computer interaction and computer-mediated communication research (Evans et al., 2017; Hutchby, 2001), definitions vary. We built on scholars such as Boyd (2010), Bucher and Helmond (2017), DeVito et al. (2017), and Van Raemdonck and Pierson (2021), who have already identified various features and affordances on social media services, such as self-presentation (DeVito et al., 2017), communicative affordances (Bucher and Helmond, 2017), and persistence (Boyd, 2010). Mainly, we make use of the taxonomy by Van Raemdonck and Pierson (2021), on ”affordances that focus on norm building within the ‘networked publics’.” We use the taxonomy to categorise affordances along visibility, network interactions and intervention (Van Raemdonck and Pierson, 2021). In addition, we analysed which affordances structured the user’s self-presentation based on the framework by DeVito et al. (2017) under the category of profile work. After identifying features and categorising them into affordances, we compare affordances between groups. Using our understanding of how particular affordances describe normative relations to user behaviour, we consider them as value statements of the design groups. inline]More on how we actually did the coding work
4. Results
Our analysis identified differences in the overall system architecture, functionality around posting and showing content, and profiles. These sections are intentionally kept ‘close to the data’ and avoid drawing interpretations to allow us to demonstrate the case. The final subsection abstracts the design choices towards their affordances and brings ideological differences into the analysis.
4.1. General Architecture
(a) Group 3 news feed
(b) Group 1 forum main page
Figure 3. Main pages of two different groups
The design brief allowed for different interpretations of the goal. Groups 1, 2, and 4 presented artefacts more like a traditional Internet forum, while Group 3’s artefact mimicked social media with a news feed.
Within forums, architectures varied. Group 1 structured the discussion forum through organisational units (such as health and welfare services, organised publicly and locally in Finland), followed by sub-forums based on ongoing projects (see Figure 3(b)). Group 2, on the other hand, structured the discussions into three groups: political parties, wider civic society, and other discussion. Similarly, Group 4 had a mixed approach based on thematic and location-based structures. With the structuring of the forum, Groups 1 and 4 added elements allowing citizen initiatives on establishing new sub-forums, showing a somewhat more citizen-centric approach to participation, going beyond just administratively defined forums.
The news feed-based design (see Figure 3(a)) did not have a preset structure but used hashtag-based taxonomy to allow user-generated structuring of the content.
4.2. Content and Posting
Only groups 3 and 4 specified how new content would be submitted to these forums. Both focused on more structured input opportunities to help citizens justify their positions. Group 4 designed an interface which scaffolds the post generation by asking user to enter their opinion and justifications on separated text input, with the assumption that this would lead to a higher quality debate (see Figure 4). In addition, they asked people to provide links to support their stance and recommended news content related to the opinion. Group 3 identified the same issue but considered an artificial intelligence agent (shown as a dog in Figure 3(a)) to suggest to the user that they could provide justifications. Their system also provided users with controls for who could view or react to their content. Interestingly, Group 3 introduced a character limit, focusing on shorter messages and propagation.
Figure 4. Scaffolding posts
Beyond providing written content and comments, two of the artefacts included a mechanism to react to the content. Group 1’s artefact used stars to show support, and such reactions would increase the topic’s visibility on lists based on what they called a “Reddit approach” This means there was competition for visibility, where more starred content would gain more visibility. Group 3 provided various reaction icons and the ability to recirculate content on their artefacts, with the aim of ‘allow showing support.’ This articulation is arguably different from Group 1’s, which did not consider reactions a mechanism for gaining popularity but emphasised their meaning for socio-emotional support in the final presentation. Groups 2 and 4 did not introduce this kind of affordance to the artefacts, with Group 2 articulating how reactions, such as ‘likes‘ are not beneficial for constructive discussion and lead to attention seeking.
Recent academic discussions on social media have focused on content moderation and its challenges (Rantasila et al., 2022). Remarkably, only groups 1 and 3 considered any form of content moderation in the final artefact, even though the focus was to devise an online space for political discussion. Group 1 hints at rules for the platform, which are enforced by moderators selected by some authority, which exemplifies the intention to govern the content (Rantasila et al., 2022). While it remains unclear who is part of the moderation team, the moderation must be conducted in a team effort. Group 3 describes a more user-driven approach to moderation with options to report posts, add trigger warnings, and show reminders to enforce rules and block users. Similarly, Group 3 mentions that there is a human editor to conduct fact-checking. This shows a mix of hard (e.g. removal of reported posts) and soft (e.g. trigger warnings) moderation styles (Rantasila et al., 2022). While we can surmise rules against spam and hate speech from Group 3 suggesting these as features, no group clearly described what content is warrant or sanctioned on their sites.
4.3. Profiles
Group 1 Group 2 Group 3 Group 4
Picture
Verification of authentication
Username
Real name
Description
Age
Party affiliation
Occupation
List of political interests
Follower counts
Indicator for a parody account
Number of warnings
Table 3. Design of user profile across groups
All four designs contained the idea of a persistent user identity as a core part of the platform. At the same time, there was a lot of diversity in the details that should be included in the artefacts. All sites included some profile pictures. Group 2 allowed pseudonyms and a person to change their username, which would keep the posting history intact. Instead, groups 1, 3 and 4 enforced a real-name policy. However, Group 4 hinted at the possibility of parody accounts, although marked by an indicator. The policy of enforcing clear names is often debated also on social media services such as Facebook (Haimson and Hoffmann, 2016).
In terms of user verification, it was discussed in all final presentations. Group 1 argued that the system should have strong authentication as “the [discussion] data is used for political decision-making” but it also was an approach to ensure proper behaviour on the platform, and Group 3 argued in the same line that there should be a mechanism to ensure that the poster ‘is a real person’ and not pretend to be some other person. Similar arguments were also presented in Group 2 and Group 4.
Interestingly, the differences do not stop here (see Table 3). Profiles included a free-form description field in groups 3 and 4. Group 2 did not design such an element; instead, after the user name, an indication of the user’s party affiliation was shown. Group 4 added age and occupation in addition to the free-formed field, as well as a list of users’ political interests. Group 3 instead included, for example, the number of warnings the user had received for breaking any rules.
4.4. From Features to Affordances to Ideologies
The following section describes the differences and similarities of the artefacts along four key affordances (each row in Table 4). We identified the key affordances popularity-based visibility, open network interactions, user intervention and flexible profile work through categorising the features mentioned above based on current understanding of social media affordances (Boyd, 2010; Van Raemdonck and Pierson, 2021; DeVito et al., 2017). In the following paragraphs, we will go through each key affordance and the differences and commonalities between groups, along with their ideological leanings.
Affordances Group 1 Group 2 Group 3 Group 4
Right-leaning Left-leaning Left-leaning Mixed
Popularity-based Visibility Encourage Refuse Discourage
Open Network Interactions Refuse Allow Encourage Refuse
User Intervention Allow Encourage Encourage
Flexible Profile Work Discourage Allow Encourage Allow
Table 4. Key Affordances of the design artefacts of each group (including their ideological composition).
According to Davis and Chouinard (2016): Encourage prompts specific actions while suppressing others; Allow remains neutral, indifferent to the action’s usage and outcome; Discourage makes an action more difficult to achieve; Refuse blocks certain actions outright.
The groups varied in their stance towards the mechanics of content visibility on their platforms: both the right-leaning and mixed composition groups preferred popularity-based metrics to promote content on the platform. Posts would have a like count, prioritising their visibility on the platform. In Group 1, popular content was prominently displayed with stars on the front page. Therefore, both groups encouraged popularity-based visibility by fostering a usage pattern that aims for posts that receive more reactions and, therefore, get rewarded with more visibility, similar to platforms like Facebook (Van Raemdonck and Pierson, 2021). Both left-leaning Groups 2 and 3 rejected popularity-based design approaches. Strikingly, as we observed above, Group 3 mentions the need for reactions and likes, but they highlight them as easy ways to show support, not as a popularity-based logic. Therefore, Group 3 considered discouraging behaviour that rewards the popularity of content with higher visibility on the platform, whereas Group 2 outright refused it by not offering any way for users to engage with content based on its popularity.
Regarding interaction forms, all but Group 3 opted for a subforum architecture. Group 1, 2 and 4 all decided to section the discussion on their platform according to organisational levels (Group 1); parties, society and others (Group 2); and themes or location (Group 4). Whether a platform technically allows anyone to see and interact with every user’s content, there is still a clear difference between a system requesting users to post in particular subforums and one accumulating posts on a news feed. The former can be considered closed group interactions because they happen in clearly delineated spaces, and content does not automatically reach users across a platform (Van Raemdonck and Pierson, 2021). Such architecture entrusts the site administrators and moderators responsible for sanctioning or rewarding certain behaviours. Contrary to that, the affordances of Group 3’s artefact enabled open network interactions, as content moves freely across the platform. Such an architecture focuses on the sender instead of an overarching group space like a subforum. This puts more responsibility and power onto the senders, as they have more possibilities for intervention and structuring the debate (Van Raemdonck and Pierson, 2021). This shows a contrast to the other left-leaning Group 2, as they chose the more conservative forum solution for their platform architecture.
Regarding interventions, we observed various approaches and remarkable differences in the way the Groups’ artefact afforded users to be involved. The right-leaning Group 1 considered the moderation rules so important that one of their artefacts was a list of written rules. While their system imposed a governing style to moderation (Rantasila et al., 2022), they still allowed user interventions through the option to report to content and initiatives for new subforums, which shows a neutral stance. The left-leaning Group 3 instead used self-guided moderation strategies with trigger warnings and warning counts (which could lead to banning from the platform) (Rantasila et al., 2022). Further, together with the mixed Group 4, they thought of structured content input that promotes user in adding justification for their posts. Thus, both groups’ artefacts encourage intervention through users.
Finally, it seems that groups including left-leaning participants provided more flexibility for users to present themselves, encouraging or at least not restricting flexible profile work (DeVito et al., 2017). While all groups required a persistent identity from their users, features such as a changeable username (Group 2) or parody accounts (Group 3) would encourage users to express themselves freely and differently. Additionally, Group 3 offered users control over the visibility of their content, encouraging users to perform self-presentation by catering to different audiences (DeVito et al., 2017). In stark contrast, the right-leaning group offered the least opportunities to customise profiles or present freely, which indicates an expectation for the users to present persistent roles on their platform and an affordance that discourages flexible profile work.
Our study only had one right-leaning and one mixed composition group due to participant recruitment challenges (see foonote 1). However, two left-leaning groups already allow us to highlight diversity within the same end of the political spectrum, highlighting how various other factors may also impact design work. For example, with profile work, Group 3 was much more creative in terms of attributes (see Table 3). Further, Group 2 subscribes to the forum approach, which separates interactions on the platforms in pre-defined groups and gives less intervention capabilities to individual users.
5. Discussion
Our findings continue to present the relevance of political ideologies as value frameworks in the context of technology design. We observed that affordances were used diversely between the particular leaning groups; we also observed that the mixed group seemed to adapt ideas that were present in right-leaning and left-leaning groups, speaking on their ability to develop ideas across sectors.
5.1. Design by the Left and the Right
Political science has various characterisations of the left and the right: citizens and experts identify this continuum and can position themselves into it (see e.g., Cochrane, 2015b). Using political party manifestos, scholars such as Cochrane (2015a), Jahn (2011) and Däubler and Benoit (2022) have identified characteristics of right-leaning parties to have – at least a subset of – ideals such as less state intervention or support for more expansive welfare programs characterise this tension (see Table 5, first row). As these are from political party manifestos, these do not directly map into the design of digital services but rather broadly characterise the division between the left and the right around the economy and the state.
The left The right
Political party manifestos the need for fair treatment of all people, special protection for underprivileged, and end of discrimination; decreasing military expenditures and disarmament; preservation of natural resources against selfish interests and other green policy ideals, support for welfare state; need to spend money on museums, art galleries, and other cultural and leisure facilities; favourable reference to underprivileged minorities who are defined neither in economic nor in demographic terms; support for market regulation and direct government control of economy, as well as state intervention into the economic system. support for free enterprises and trust on market-based dynamics overall as well as personal initiative as well as support for traditional economic orthodoxy and a need for efficiency and economy in government and administration; support on economic incentives to boost enterprise and entrepreneurship as well as support for free trade; reduction of state-sponsored social service or social security scheme and limiting spending on education; support for traditional moral values, such as prohibition, censorship, and suppression of immorality and unseemly behaviour; less support for labour groups, working class, unemployed; enforcement or encouragement of cultural integration
Digital design principles Paternalism Rich identity and profile work Freedom and responsibility Competitive logic
Table 5. Social media design principles for the left–right axis
Moving from political manifestos to digital artifacts and design principles in them (see Table 5, second row row). Grön and Nelimarkka (2020) suggested that the left–right dimension appeared as a tension of paternalims and freedom. More extensively, they showed acceptable design approaches seemed to be rooted in the values of each party, for example, communality was preferred by a liberal-conservative and agrarian Centre Party.
We observe a similar ideological division in the role of market dynamics in controlling visibility. The right-leaning group suggested affordances encouraging popularity as a driver for content visibility, with an underlying idea of the platform as a marketplace for ideas competing for attention. Both left-leaning groups discouraged such an approach. Although traditional political party manifestos have focused on the contest within the state and markets, our work shows that concerns related to marked dynamics generalise to a broader design principle on the role of any kind of competition logistic in social media services.
Similarly, we observed differences in the formulation of user identity. Both left-leaning groups appeared to allow and encourage more flexible profile work, while the right-leaning group limited the options for users to present themselves. In part, this difference may speak to different conceptualisations of how people form social groups. For example, occupation as a primary social group is important for left-wing parties, which historically built on social class and related occupational status, know as class-mass parties (Gunther and Diamond, 2003). Therefore, it is not surprising that these fields for self-expression appeared more extensive. In addition, supporters of left-leaning parties tent to have more extraverted personality and be more open to experiences (Osborne et al., 2021). Therefore, it reasons to argue that they may embrace a more nuanced representation of identity.
5.2. Beyond Left and Right on the Ideological Spectrum
As highlighted in Section 2.1, left–right is only one ideology dimension. Our analysis and outcomes directly relate to our participant selection process, where we sought groups along this axis. We also collected users’ ideological positions on the liberal–conservative axis. Group 3 stands out as the only one with a pronounced liberal-leaning ideological measurement (see Table 1). Other groups had only a slight leaning towards the liberal end but had a high variance within the group of participants, indicating the groups were more heterogeneous on this axis.
That might explain why Group 3 was the only group that encouraged the idea of an algorithmically and profile-based structured social network and open network interactions. This form differs from the more traditional approach of structuring and interaction and might even signify an underlying agreement of slightly more conservative group identities (Lampinen et al., 2014). This points to a disparity between the two left-leaning groups and highlights the importance of investigating the influence of political ideologies in the design process of digital technologies beyond the left-right spectrum.
5.3. Limitations and Future Work
There are several limitations to our study: most vitally, our sample is small, and there is a lack of right-leaning and mixed groups, inhibiting a more extensive within-ideology comparison. We re-iterate that this was due to the challenge of reaching out to a particular political party, thus limiting our access to right-leaning participation. Nonetheless, the small sample may limit the generalisability of our findings. Similarly, we acknowledge that some results might be due to the difference in the groups’ social networking experiences, age, or other compounding factors. However, as we connect our findings to political ideologies, the results have a stronger foundation on political science than just providing empirical observations.
Second, political ideologies are highly contextualised. The outcomes may be different in another country or may change temporarily as political ideologies and their thinking evolve (for summary, see Cochrane, 2015b). For example, Nelimarkka et al. (2019) show that acceptable polarisation mitigation strategies for social media differ in Finland and the United States. These two issues call for larger studies and comparative approaches to examine how the political ideologies shape potential design language.
Finally, this analysis focused only on final artifacts and their presentation. Further research could examine more group dynamics during the group design stages. It would be possible to analysis to examine who introduced design ideas, how those were discussed and if those discussions lead to their incorporation in the final artifact. This would provide more nuanced perspectives on where the design ideas emerge and how they are examined in the group, thus also addressing power stances.
5.4. Implications to Practice
Our work emphasizes the importance of political ideologies as coherent value sets in the design of social media services. We begin to build a connection between the state – represented by its democratically elected political representatives affiliated with political parties – and technology design beyond mere regulation. Traditionally, political action in the technology domain has focused on regulating the relationship between users and platforms, such as through GDPR or the AI Act, and on regulating interactions between platform companies, such as in antitrust cases. However, we propose that there is a more general design language that outlines acceptable affordances in technology design per political ideology.
Therefore, researchers using value-based approaches or seeking to introduce ideas of ‘democratic design’ should ensure that diversity on the ideological spectrum is maintained to avoid political leanings in such works, especially if touching on applications with societal relevance. Furthermore, our work diversifies the focus on ideologies, whereas current critical scholarship focuses mainly on capitalism and feminism (Keyes et al., 2019; Bardzell, 2010). Our work highlights the diversity of ideologies and their utility in examining values.
We would not advocate a generic social media service to examine ideological diversity, as they operate across various settings. (However, we do consider it a relevant question which ideological design principles are in use in these services, cf., Kender and Frauenberger, 2022; Hasinoff and Bivens, 2021) However, for projects planning to expand social media to serve as a better avenue for political discussion, political ideals of participation are essential – concepts such as good democracy are ultimately normative (Held, 2006). Such efforts may forget how technologies manifest values, and choices such as self-presentation or competition for visibility are entangled with political values. In such cases, deliberative processes where a service is developed in a mixed ideology setting are vital to ensure the acceptability of the service across the ideological spectrum. Our mixed composition group already seemed to integrate both more left-leaning and right-leaning principles into their artifact.
6. Conclusions: Towards Political HCI?
Our study explored the design of social media services within groups formed by politically aligned elites. By examining the artifacts produced, we observed that these designs reflect ideological alignments: certain features are preferred in groups with right-leaning members, while others are favored by left-leaning members. Our findings underscore that political ideologies represent a coherent set of values even in the design domain of digital systems, extending and generalizing existing political science research on ideologies.
Political power increasingly influences the development of social computing systems through legislation and government funding for technology development. Beyond this direct influence, political ideologies manifest as broader sets of values, raising the challenge of ensuring more democratic representation of these values during system development. Therefore, social computing researchers must engage with political science and the study of political ideologies. This engagement will help understand, unpack, and make visible the political power embedded in these systems, fostering public debate and further examining digital systems’ societal impact.